A firm has weekly revenue of $1000.

a) The firm's total cost is $1450 per week. The firm will shut down if weekly fixed (sunk) cost is what?

b) The firm's variable cost is $500 and its fixed cost is $800. The firm shuts down if the percentage of fixed costs that is avoidable is greater than

20%, 40%, 60%, or 80%?
Answer #1

Revenue = $1,000

a) Firm shut down when revenue < variable cost or the firm of the price of the product is less than marginal cost.

Total cost = $1,450

Total cost = Variable cost + Fixed cost

If fixed cost is lets say 449 or less, then variable cost would be 1,001 or more which creates a situation of shutting down.

b) Variable cost = $500

Fixed cost = $800

Firm shut down when cost becomes greater than 1,000.

If 20% of the cost is avoidable while 80% is non avoidable, it means that producer will bear 80% of 800 = 640 of fixed cost in any case possible plus variable cost which turns out to be 500 + 640 = 1,140

If 40% of the cost is avoidable while 60% is non avoidable, it means that producer will bear 60% of 800 = 480 of fixed cost in any case possible plus variable cost which turns out to be 500 + 480 = 980

If 60% of the cost is avoidable while 40% is non avoidable, it means that producer will bear 40% of 800 = 320 of fixed cost in any case possible plus variable cost which turns out to be 500 + 320 = 820

If avoidable cost is greater than 20%, firm cost is more than revenue and falling. If it is above revenue, firm will shut down.